OpenStudy (anonymous):

HELP PLEASE?!?!?! Express the following function, f(x) as a composition of two functions f and g.

OpenStudy (anonymous):

\[f(x) = \frac{ x^2 }{( x^2 + 4) }\]

OpenStudy (anonymous):

well you can use reciprocal of x^2+4 do you know how?

OpenStudy (anonymous):

no normally i would set the binomial = to 0

OpenStudy (anonymous):

ok what if i told you your function looks like this a/b is same as a*1/b reciprocal just means flip it and put a 1 on top. 1/b is reciprocal.

OpenStudy (anonymous):

now for the pattern a/b=a*1/b you plug in your numbers x^2/(x^2+4)= x^2* (1/ (x^2+4) )

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well it ask to compose it of 2 functions so f(x)=x^2 and g(x)=(1/ (x^2+4) ) therefore f(x)*g(x)= x^2/(x^2+4) Unless i could be wrong and they mean f(g(x))

OpenStudy (anonymous):

well i think my teacher would take anything, because he didn't say fog or gof

OpenStudy (anonymous):

composite means f o g (x)

OpenStudy (anonymous):

that is easy too tho f(g(x)) let me try it

OpenStudy (anonymous):

you got f(x) = x/(x+4) then g(x)=x^2 then f o g (x)= f(g(x)) = x^2/(x^2+4) Cuzz you plug in x^2 into ever x in in the f(x) function f(x) = x/(x+4) ^ ^ plug in x^2